OpenStudy (anonymous):

There are curves y=x^2 and y=mx, where m is some positive constant. No matter what positive constant m is, the two curves enclose a region in the first quadrant. Find the positive constant m such that the area of the region bounded by the curves y=x^2 and y=mx is equal to 8.

OpenStudy (anonymous):

\[ \int_0^m \left(m x-x^2\right) \, dx=\frac{m^3}{6}=8 \] Solve for m

OpenStudy (anonymous):

\[ m=2 \sqrt[3]{6} \]

OpenStudy (campbell_st):

well find the points of intersection of the curves... so equate the curves \[x^2 = mx ... or ....x^2 - mx = 0\] then \[x(x - m) = 0\] so the 2 curves intesect at x = 0 and x = m here is the picture  so you need to find the area between the curves and you know the area is 8 so integrating its \[A = \int\limits_{0}^{m}mx...dx - \int\limits_{0}^{m} x^2... dx\] evaluate the integral then solve for m. hope it helps